Circular No. 6463

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

PROBABLE NOVA IN CRUX W. Liller, Viña del Mar, Chile, reports his discovery with PROBLI-

COM of a probable nova on a photograph taken on Aug. 26.04 UT. The candidate is located at α = 12^{h}10^{m}.51, δ = −61^{o}45′.3 (equinox 2000.0). Magnitude estimates: Aug. 7.0, [11.5: (Liller); 26.98, 9.25 (Liller; CCD, broad-band V system, compared to nearby GSC stars); 27.781, 10.1 (D. Overbeek, Edenvale, South Africa; visual; communicated by J. Mattei).

BL LACERTAE A. Robinson, Division of Physics and Astronomy, University of Hert-

fordshire, Hatfield, writes, with regard to IAUC 6457: “This source was observed on Aug. 20.1 UT by E. Corbett and myself using the ISIS spectrograph of the 4.2-m William Herschel Telescope. A broad Hα emission line is clearly present in our spectrum with equivalent width 0.5 ± 0.04 nm and FWHM ∼ 4000 km/s. Narrow [N II] 658.4-nm (FWHM ∼ 300 km/s) is also present, superimposed on the Hα line.”

COMET C/1995 O1 (HALE-BOPP) Numerous observers have reported steady, strong jetting activity from

both visual and CCD observations of this comet during the past few months. The jets, including a noticeable northward fan, have significantly increased in visibility since May. Visual observations on Aug. 14–15 by S. J. O’Meara (Volcano, HI, 0.12-m refractor) and by J. E. Bortle (Stormville, NY, 0.41m reflector) reveal a bright, “J”-shaped, ever-widening fan spanning some 40^{o}–50^{o} centered close to due north and extending up to 10′ in length. K. Birkle, Centro Astronómico Hispano-Aleman, reports that CCD exposures with the 1.2-m telescope during May 29–June 25 using various filters show the northward fan as three separate structures; a 30-s R image with the 2.2m reflector on July 19.86 UT (taken with the help of E. Thommes and H. H. Hippelein) also shows low-contrast features in p.a. 45^{o} and 70^{o} that were also possibly present in May, while two brighter structures rotated from p.a. 156^{o} to 142^{o} and from 235^{o} to 278^{o} during the 7-week interval. F. Manzini, C. Guaita and F. Crippa write that CCD images obtained with the 0.33-m reflector at the SAS Observatory, near Milan, show jets near p.a. 40^{o}, 145^{o} and 270^{o} to be very faint on May 28 but almost as bright as the main tail by July 31, with further brightening and the development of several secondary jets by Aug. 16; in view of the numerous linear spikes extending 10′′–15′′ from the nucleus and having little tendency to bend they remark that the comet resembles a porcupine.

1996 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden